# Location–Scale Calibration for Generalized Posterior

### v2

@@ Appendix D Non-smooth penalties: active set and subgradient calculus
$\boldsymbol{0}\in\Psi(\boldsymbol{\theta})+\lambda\partial\rho(\boldsymbol{\theta})$
with solution $\boldsymbol{\theta}^{\lambda}$, where $\partial\rho$ is the convex subdifferential. For an index set $A\subset\{1,\dots,p\}$, we write $\boldsymbol{\theta}_{A}:=(\theta_{j})_{j\in A}$ for the subvector of $\boldsymbol{\theta}$ with coordinates in $A$, and, for any matrix $\boldsymbol{M}$, we write $\boldsymbol{M}_{AA}:=P_{A}\boldsymbol{M}P_{A}^{\top}$ for the corresponding principal submatrix, where $P_{A}$ denotes the coordinate projection onto $A$.
@@
The penalty $\rho:\mathbb{R}^{p}\to(-\infty,\infty]$ is convex and lower semicontinuous. The penalized population inclusion has a unique solution $\boldsymbol{\theta}^{\lambda}$. Moreover, there exists an index set
$A:=\bigl\{j:\theta^{\lambda}_{j}\neq 0\bigr\}\cup\bigl\{j:0\in\partial\rho_{j}(\boldsymbol{\theta}^{\lambda})\bigr\}$
and a neighborhood $\mathfrak{N}_{A}$ of $\boldsymbol{\theta}^{\lambda}_{A}$ such that the map $\boldsymbol{\theta}_{A}\mapsto\rho(\boldsymbol{\theta}_{A},\boldsymbol{\theta}_{A^{c}}^{\lambda})$ is twice continuously differentiable on $\mathfrak{N}_{A}$; the Clarke generalized Jacobian $\partial^{\mathrm{G}}\nabla\rho(\boldsymbol{\theta}^{\lambda})$ is nonempty and contains a symmetric positive semidefinite matrix $\boldsymbol{H}_{\rho}(\boldsymbol{\theta}^{\lambda})$; and the matrix $\boldsymbol{J}^{\star}_{\lambda}:=\boldsymbol{J}^{\star}+\lambda\boldsymbol{H}_{\rho}(\boldsymbol{\theta}^{\lambda})$ is nonsingular, where $\boldsymbol{J}^{\star}:=-\nabla_{\boldsymbol{\theta}}\Psi(\boldsymbol{\theta}^{\lambda})$ and $\boldsymbol{K}^{\star}:=\operatorname{Var}_{P^{\star}}[\psi(\mathcal{D}_{1},\boldsymbol{\theta}^{\lambda})]$.
Throughout this section we work on the active set $A$ supplied by Assumption D.1. Under Assumption D.1, we define
$\boldsymbol{V}_{\mathrm{target}}^{\star}:=\bigl(\boldsymbol{J}_{\lambda}^{\star}\bigr)^{-1}\boldsymbol{K}^{\star}\bigl(\boldsymbol{J}_{\lambda}^{\star}\bigr)^{-1},$
exactly as in the smooth case, and we denote by $\boldsymbol{J}^{\star}_{AA}$, $\boldsymbol{K}^{\star}_{AA}$, and so on, the corresponding $AA$-blocks.
Lemma D.2 (Active-set smoothness and curvature).
Suppose Assumption D.1 holds and define
@@
$\boldsymbol{H}_{\rho,AA}:=\left.\nabla^{2}_{\boldsymbol{\theta}_{A}}\rho\bigl(\boldsymbol{\theta}_{A},\boldsymbol{\theta}^{\lambda}_{A^{c}}\bigr)\right|_{\boldsymbol{\theta}_{A}=\boldsymbol{\theta}^{\lambda}_{A}}.$
Then $\boldsymbol{H}_{\rho,AA}$ exists, is symmetric positive semidefinite, and coincides with the $AA$-block of any symmetric selection $\boldsymbol{H}_{\rho}(\boldsymbol{\theta}^{\lambda})\in\partial^{\mathrm{G}}\nabla\rho(\boldsymbol{\theta}^{\lambda})$, that is,
$\boldsymbol{H}_{\rho,AA}=P_{A}\boldsymbol{H}_{\rho}(\boldsymbol{\theta}^{\lambda})P_{A}^{\top}.$
Consequently,
$\boldsymbol{J}_{\lambda,AA}^{\star}:=\boldsymbol{J}^{\star}_{AA}+\lambda\boldsymbol{H}_{\rho,AA}$
is nonsingular whenever $\boldsymbol{J}_{\lambda}^{\star}$ is nonsingular. For the $\ell_{1}$-penalty $\rho(\boldsymbol{\theta})=\sum_{j=1}^{p}|\theta_{j}|$, we have $\boldsymbol{H}_{\rho,AA}=\boldsymbol{0}$.
Proof of Lemma D.2.
By Assumption D.1, $\rho$ is convex and lower semicontinuous and is $C^{2}$ in the active coordinates on a neighborhood of $\boldsymbol{\theta}^{\lambda}_{A}$, which gives the existence and positive semidefiniteness of $\boldsymbol{H}_{\rho,AA}$. The Clarke generalized Jacobian $\partial^{\mathrm{G}}\nabla\rho(\boldsymbol{\theta}^{\lambda})$ is the convex hull of limits of classical Hessians at differentiability points approaching $\boldsymbol{\theta}^{\lambda}$. Since $\rho$ is $C^{2}$ in the active coordinates in a neighborhood of $\boldsymbol{\theta}_{A}^{\lambda}$, every such limit has the same $AA$-block, which must therefore coincide with $\boldsymbol{H}_{\rho,AA}$ for any symmetric selection $\boldsymbol{H}_{\rho}(\boldsymbol{\theta}^{\lambda})$. The nonsingularity of $\boldsymbol{J}_{\lambda,AA}^{\star}$ follows from that of $\boldsymbol{J}_{\lambda}^{\star}$ and the block structure.
For the $\ell_{1}$-penalty, each coordinate map $\theta_{j}\mapsto|\theta_{j}|$ is affine on any interval not containing the origin. Under sign/gap stability, $\theta^{\lambda}_{j}\neq 0$ for $j\in A$ and the neighborhood of $\boldsymbol{\theta}^{\lambda}_{A}$ can be chosen so that $\operatorname{sgn}(\theta_{j})$ is constant for $j\in A$. Hence $\rho$ is locally linear in $\boldsymbol{\theta}_{A}$ and the second derivative vanishes, giving $\boldsymbol{H}_{\rho,AA}=\boldsymbol{0}$. ∎
@@
such that $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\to_{p}\boldsymbol{\theta}^{\lambda}$ and
$\Pr\!\left(\operatorname{sgn}\bigl(\hat{\theta}^{\mathrm{pen}}_{n,j}\bigr)=\operatorname{sgn}\bigl(\theta^{\lambda}_{j}\bigr)\text{ for all }j\in A\right)\longrightarrow 1.$
The latter is the usual sign/gap stability requirement for the active set.
@@
Suppose Assumptions D.1, 2.2 and 2.3 hold, and let $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}$ satisfy (23) with the above sign/gap stability. Then
$\sqrt{s_{n}}\bigl(P_{A}\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}-P_{A}\boldsymbol{\theta}^{\lambda}\bigr)\longrightarrow_{d}\mathcal{N}\Bigl(\boldsymbol{0},\bigl(\boldsymbol{J}_{\lambda,AA}^{\star}\bigr)^{-1}\boldsymbol{K}^{\star}_{AA}\bigl(\boldsymbol{J}_{\lambda,AA}^{\star}\bigr)^{-1}\Bigr),\quad n\to\infty.$
@@
On the event where $\operatorname{sgn}(\hat{\theta}^{\mathrm{pen}}_{n,j})=\operatorname{sgn}(\theta^{\lambda}_{j})$ for all $j\in A$ and $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}$ lies in the neighborhood $\mathfrak{N}_{A}$ from Lemma D.2, the restriction of $\rho$ to the active coordinates is $C^{2}$ and the subgradient in those coordinates is a singleton: there exists a measurable choice $\nabla_{A}\rho(\boldsymbol{\theta})$ such that
$\partial\rho(\boldsymbol{\theta})\cap\mathbb{R}^{A}=\left\{\nabla_{A}\rho(\boldsymbol{\theta})\right\},\quad\boldsymbol{\theta}_{A}\in\mathfrak{N}_{A}.$
Projecting (23) onto $A$ and using this selection gives the smooth estimating equation
$\boldsymbol{0}=P_{A}\boldsymbol{U}_{n}\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\bigr)-\lambda_{n}\nabla_{A}\rho\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\bigr).$
@@
A first-order Taylor expansion of both $\boldsymbol{U}_{n}$ and $\nabla_{A}\rho$ at $\boldsymbol{\theta}^{\lambda}$ yields
@@
$\boldsymbol{0}=P_{A}\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\lambda_{n}\nabla_{A}\rho\bigl(\boldsymbol{\theta}^{\lambda}\bigr)+\boldsymbol{J}^{\star}_{AA}\bigl(\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}\bigr)+\lambda_{n}\boldsymbol{H}_{\rho,AA}\bigl(\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}\bigr)+o_{p}\!\left(\bigl\|\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}\bigr\|\right)$
on $A$, uniformly on events of probability tending to one. Using the population inclusion $\boldsymbol{0}\in\Psi(\boldsymbol{\theta}^{\lambda})-\lambda\partial\rho(\boldsymbol{\theta}^{\lambda})$ and the definition of $\boldsymbol{J}^{\star}$, this can be written as
@@
$\boldsymbol{0}=P_{A}\bigl\{\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\bigr\}+\boldsymbol{J}_{\lambda,AA}^{\star}\bigl(\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}\bigr)+o_{p}\!\left(\bigl\|\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}\bigr\|\right).$
@@
$\sqrt{s_{n}}P_{A}\bigl\{\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\bigr\}\longrightarrow_{d}\mathcal{N}\bigl(\boldsymbol{0},\boldsymbol{K}^{\star}_{AA}\bigr).$
Since $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\to_{p}\boldsymbol{\theta}^{\lambda}$ and $\boldsymbol{J}_{\lambda,AA}^{\star}$ is nonsingular by Lemma D.2, the remainder term is $o_{p}(s_{n}^{-1/2})$ and the usual $M$-estimation argument yields
$\sqrt{s_{n}}\bigl(\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}\bigr)=-\bigl(\boldsymbol{J}_{\lambda,AA}^{\star}\bigr)^{-1}\sqrt{s_{n}}P_{A}\bigl\{\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\bigr\}+o_{p}(1).$
The desired normal limit follows by Slutsky’s theorem. ∎
Lemmas D.2 and D.3 show that, under the non-smooth setting, the active coordinates behave as in the smooth case, with curvature matrix $\boldsymbol{J}_{\lambda,AA}^{\star}$ and variability $\boldsymbol{K}_{AA}^{\star}$. The proofs of Proposition 3.1, Theorem 3.2, Lemma 4.1, Lemma 4.2, and Proposition 4.3, use only local quadratic expansions and the central limit theorem. Repeating those arguments with all matrices and vectors restricted to $A$ yields the same Gaussian limits on the active coordinates after replacing $\boldsymbol{J}^{\star}$, $\boldsymbol{J}_{\lambda}^{\star}$, $\boldsymbol{K}^{\star}$, $\boldsymbol{V}_{\mathrm{target}}^{\star}$ by their $AA$-blocks. Coordinates in $A^{c}$ may exhibit boundary phenomena, and their asymptotic distribution need not be Gaussian. Accordingly, our asymptotic statements and calibrated inference for non-smooth penalties are reported conditionally on the active set $A$.

### v3

@@ Appendix D Non-smooth penalties: active set and subgradient calculus
$\boldsymbol{0}\in\Psi(\boldsymbol{\theta})+\lambda\partial\rho(\boldsymbol{\theta})$ (23)
with solution $\boldsymbol{\theta}^{\lambda}$, where $\partial\rho$ is the convex subdifferential. For an index set $A\subset\{1,\ldots,p\}$, we write $\boldsymbol{\theta}_{A}:=(\theta_{j})_{j\in A}$ for the subvector of $\boldsymbol{\theta}$ with coordinates in $A$, and, for any matrix $\boldsymbol{M}$, we write $\boldsymbol{M}_{AA}:=P_{A}\boldsymbol{M}P_{A}^{\top}$ for the corresponding principal submatrix, where $P_{A}$ denotes the coordinate projection onto $A$. For any function $g:\mathbb{R}^{p}\to\mathbb{R}$ that is differentiable at $\boldsymbol{\theta}$, we define the active gradient
$\nabla_{A}g(\boldsymbol{\theta}):=P_{A}\nabla g(\boldsymbol{\theta})\in\mathbb{R}^{|A|}$
and, when $g$ is twice differentiable at $\boldsymbol{\theta}$, the active Hessian
$\nabla^{2}_{\boldsymbol{\theta}_{A}}g(\boldsymbol{\theta}):=P_{A}\nabla^{2}g(\boldsymbol{\theta})P_{A}^{\top}\in\mathbb{R}^{|A|\times|A|}.$
Whenever we write $\nabla_{A}\rho(\boldsymbol{\theta})$ or $\nabla^{2}_{\boldsymbol{\theta}_{A}}\rho(\boldsymbol{\theta})$, we implicitly restrict attention to points $\boldsymbol{\theta}$ at which these quantities are well defined.
@@
The penalty $\rho:\mathbb{R}^{p}\to(-\infty,\infty]$ is convex and lower semicontinuous. The penalized population inclusion (23) has a unique solution $\boldsymbol{\theta}^{\lambda}$. Moreover, there exists an index set $A\subset\{1,\ldots,p\}$ and an open neighborhood $\mathcal{N}$ of $\boldsymbol{\theta}^{\lambda}$ such that, for every $\boldsymbol{\theta}=(\boldsymbol{\theta}_{A},\boldsymbol{\theta}_{A^{c}})\in\mathcal{N}$, the map $\boldsymbol{\theta}_{A}\mapsto\rho(\boldsymbol{\theta}_{A},\boldsymbol{\theta}_{A^{c}})$ is twice continuously differentiable on a neighborhood of $\boldsymbol{\theta}^{\lambda}_{A}$, and the Hessian with respect to $\boldsymbol{\theta}_{A}$, $\nabla^{2}_{\boldsymbol{\theta}_{A}}\rho(\boldsymbol{\theta}_{A},\boldsymbol{\theta}_{A^{c}})$, is continuous at $\boldsymbol{\theta}^{\lambda}$. Finally, the active-set penalized curvature matrix $\boldsymbol{J}_{\lambda,AA}^{\star}$ defined below is nonsingular.
We refer to $A$ in Assumption D.1 as the active set. By convexity and the above partial $C^{2}$-smoothness in the active coordinates, the Hessian
@@
$\boldsymbol{H}_{\rho,AA}:=\left.\nabla^{2}_{\boldsymbol{\theta}_{A}}\rho\bigl(\boldsymbol{\theta}_{A},\boldsymbol{\theta}_{A^{c}}\bigr)\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\lambda}}$
exists and is symmetric positive semidefinite. Moreover, shrinking $\mathcal{N}$ if necessary, we may assume that the partial gradient $\nabla_{A}\rho(\boldsymbol{\theta})$ is well defined for all $\boldsymbol{\theta}\in\mathcal{N}$ and continuous on $\mathcal{N}$, and that, for each fixed $\boldsymbol{\theta}_{A^{c}}$ with $(\boldsymbol{\theta}_{A}^{\lambda},\boldsymbol{\theta}_{A^{c}})\in\mathcal{N}$, the map $\boldsymbol{\theta}_{A}\mapsto\nabla_{A}\rho(\boldsymbol{\theta}_{A},\boldsymbol{\theta}_{A^{c}})$ is continuously differentiable at $\boldsymbol{\theta}_{A}^{\lambda}$ with derivative $\boldsymbol{H}_{\rho,AA}$. Recalling
$\boldsymbol{J}^{\star}:=\nabla_{\boldsymbol{\theta}}\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr),\quad\boldsymbol{K}^{\star}:=\operatorname{var}_{P^{\star}}\bigl(\psi(\mathcal{D}_{1},\boldsymbol{\theta}^{\lambda})\bigr),$
we define the active-set penalized curvature
$\boldsymbol{J}_{\lambda,AA}^{\star}:=\boldsymbol{J}^{\star}_{AA}+\lambda\boldsymbol{H}_{\rho,AA}.$
We also set
$\boldsymbol{V}_{\mathrm{target},AA}^{\star}:=\bigl(\boldsymbol{J}_{\lambda,AA}^{\star}\bigr)^{-1}\boldsymbol{K}^{\star}_{AA}\bigl(\boldsymbol{J}_{\lambda,AA}^{\star}\bigr)^{-1},$
which is the analogue of $\boldsymbol{V}_{\mathrm{target}}^{\star}$ in the smooth case, restricted to the active set.
Example D.2 ($\ell_{1}$ penalty).
Consider the $\ell_{1}$ penalty $\rho(\boldsymbol{\theta})=\sum_{j=1}^{p}|\theta_{j}|$. A natural choice of active set is
$A:=\{j:\theta_{j}^{\lambda}\neq 0\}.$
For each $j\in A$ we have $\theta_{j}^{\lambda}\neq 0$, so the map $x\mapsto|x|$ is affine in a neighborhood of $\theta_{j}^{\lambda}$. Hence
$\boldsymbol{\theta}_{A}\mapsto\rho\bigl(\boldsymbol{\theta}_{A},\boldsymbol{\theta}^{\lambda}_{A^{c}}\bigr)=\sum_{j\in A}|\theta_{j}|+\sum_{j\notin A}|\theta_{j}^{\lambda}|$
is affine in $\boldsymbol{\theta}_{A}$ on a neighborhood of $\boldsymbol{\theta}^{\lambda}_{A}$, and therefore twice continuously differentiable there with Hessian $\boldsymbol{H}_{\rho,AA}=\boldsymbol{0}$. Thus Assumption D.1 is satisfied with this choice of $A$ and $\boldsymbol{H}_{\rho,AA}=\boldsymbol{0}$.
@@
such that $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\to_{p}\boldsymbol{\theta}^{\lambda}$.
@@
Suppose Assumptions D.1 and 2.2 hold, and let $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}$ satisfy (24) with $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\to_{p}\boldsymbol{\theta}^{\lambda}$ and $\lambda_{n}\to\lambda$ with $s_{n}^{1/2}(\lambda_{n}-\lambda)\to 0$.
$s_{n}^{1/2}\bigl\|P_{A^{c}}\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}-P_{A^{c}}\boldsymbol{\theta}^{\lambda}\bigr\|\longrightarrow_{p}0.$ (25)
Then
$s_{n}^{1/2}\bigl(P_{A}\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}-P_{A}\boldsymbol{\theta}^{\lambda}\bigr)\longrightarrow_{d}N\Bigl(\boldsymbol{0},\bigl(\boldsymbol{J}_{\lambda,AA}^{\star}\bigr)^{-1}\boldsymbol{K}^{\star}_{AA}\bigl(\boldsymbol{J}_{\lambda,AA}^{\star}\bigr)^{-1}\Bigr),\quad n\to\infty.$
@@
By Assumption D.1 and the discussion following it, shrinking $\mathcal{N}$ if necessary, we may assume that the partial gradient $\nabla_{A}\rho(\boldsymbol{\theta})$ is well defined and continuous for all $\boldsymbol{\theta}\in\mathcal{N}$, and that, for each fixed $\boldsymbol{\theta}_{A^{c}}$ with $(\boldsymbol{\theta}_{A}^{\lambda},\boldsymbol{\theta}_{A^{c}})\in\mathcal{N}$, the map $\boldsymbol{\theta}_{A}\mapsto\nabla_{A}\rho(\boldsymbol{\theta}_{A},\boldsymbol{\theta}_{A^{c}})$ is continuously differentiable at $\boldsymbol{\theta}_{A}^{\lambda}$ with derivative $\boldsymbol{H}_{\rho,AA}$. Since $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\to_{p}\boldsymbol{\theta}^{\lambda}$, we may and do work on the event $\{\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\in\mathcal{N}\}$, whose probability tends to one. On this event $\rho$ is differentiable in the active coordinates at $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}$, so by standard convex analysis the projection of any subgradient onto the active coordinates is unique: for every $\boldsymbol{\zeta}\in\partial\rho(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}})$ we have
$P_{A}\boldsymbol{\zeta}=\nabla_{A}\rho\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\bigr).$
Since $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}$ solves (24), there exists a measurable choice $\boldsymbol{\zeta}_{n}\in\partial\rho(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}})$ such that
$\boldsymbol{0}=\boldsymbol{U}_{n}\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\bigr)+\lambda_{n}\boldsymbol{\zeta}_{n}.$
Projecting onto the active coordinates and using the above identification gives
$\boldsymbol{0}=P_{A}\boldsymbol{U}_{n}\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\bigr)+\lambda_{n}\nabla_{A}\rho\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\bigr).$
@@
A first-order Taylor expansion of both $\boldsymbol{U}_{n}$ and $\nabla_{A}\rho$ at $\boldsymbol{\theta}^{\lambda}$, using Assumption 2.2 (ii) – (iii) of the main text and the continuity of $\nabla^{2}_{\boldsymbol{\theta}_{A}}\rho$ from Assumption D.1, yields, on the active coordinates,
@@
$\begin{split}\boldsymbol{0}&=P_{A}\boldsymbol{U}_{n}\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\bigr)+\lambda_{n}\nabla_{A}\rho\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\bigr),\\ &=P_{A}\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)+\lambda_{n}\nabla_{A}\rho\bigl(\boldsymbol{\theta}^{\lambda}\bigr)+\boldsymbol{J}^{\star}_{AA}\bigl(\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}\bigr)+\boldsymbol{J}^{\star}_{AA^{c}}\bigl(\hat{\boldsymbol{\theta}}_{n,A^{c}}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A^{c}}\bigr)\\ &\qquad+\lambda_{n}\boldsymbol{H}_{\rho,AA}\bigl(\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}\bigr)+\boldsymbol{R}_{n},\end{split}$
where $\boldsymbol{R}_{n}$ collects the higher-order remainder terms arising from the Taylor expansions in the active coordinates and the dependence of $\nabla_{A}\rho$ on $\boldsymbol{\theta}_{A^{c}}$. By the smoothness assumptions and the consistency $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}\to_{p}\boldsymbol{\theta}^{\lambda}$, we may choose $\boldsymbol{R}_{n}$ so that
@@
$\|\boldsymbol{R}_{n}\|=o_{p}\!\left(\bigl\|\hat{\boldsymbol{\theta}}_{n}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}\bigr\|\right).$ (26)
Moreover, (25) implies
$s_{n}^{1/2}\bigl\|\boldsymbol{J}^{\star}_{AA^{c}}\bigl(\hat{\boldsymbol{\theta}}_{n,A^{c}}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A^{c}}\bigr)\bigr\|=o_{p}(1).$
In the remainder of the proof we absorb the cross term $\boldsymbol{J}^{\star}_{AA^{c}}(\hat{\boldsymbol{\theta}}_{n,A^{c}}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A^{c}})$ into $\boldsymbol{R}_{n}$, so that (27) continues to hold for the redefined remainder.
By the population inclusion (23), there exists $\boldsymbol{\zeta}^{\lambda}\in\partial\rho(\boldsymbol{\theta}^{\lambda})$ such that
$\Psi(\boldsymbol{\theta}^{\lambda})+\lambda\boldsymbol{\zeta}^{\lambda}=\boldsymbol{0}.$
Since $\rho$ is differentiable in the active coordinates at $\boldsymbol{\theta}^{\lambda}$, we also have
$P_{A}\boldsymbol{\zeta}^{\lambda}=\nabla_{A}\rho\bigl(\boldsymbol{\theta}^{\lambda}\bigr).$
Hence
$P_{A}\Psi(\boldsymbol{\theta}^{\lambda})+\lambda\nabla_{A}\rho(\boldsymbol{\theta}^{\lambda})=\boldsymbol{0}.$
Subtracting this identity from the Taylor expansion above, and collecting the terms in $\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}$, gives
$\boldsymbol{0}=P_{A}\bigl\{\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\bigr\}+(\lambda_{n}-\lambda)\nabla_{A}\rho\bigl(\boldsymbol{\theta}^{\lambda}\bigr)+\boldsymbol{J}_{\lambda_{n},AA}^{\star}\bigl(\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}\bigr)+\boldsymbol{R}_{n},$
where we have written
$\boldsymbol{J}_{\lambda_{n},AA}^{\star}:=\boldsymbol{J}^{\star}_{AA}+\lambda_{n}\boldsymbol{H}_{\rho,AA}.$
By continuity of $\lambda\mapsto\boldsymbol{J}^{\star}_{AA}+\lambda\boldsymbol{H}_{\rho,AA}$ and the nonsingularity of $\boldsymbol{J}_{\lambda,AA}^{\star}$ from Assumption D.1, $\boldsymbol{J}_{\lambda_{n},AA}^{\star}$ is nonsingular for all sufficiently large $n$, and $\boldsymbol{J}_{\lambda_{n},AA}^{\star}\to\boldsymbol{J}_{\lambda,AA}^{\star}$.
Rearranging the preceding display gives
$\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}=-\bigl(\boldsymbol{J}_{\lambda_{n},AA}^{\star}\bigr)^{-1}P_{A}\bigl\{\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\bigr\}-(\lambda_{n}-\lambda)\bigl(\boldsymbol{J}_{\lambda_{n},AA}^{\star}\bigr)^{-1}\nabla_{A}\rho\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\bigl(\boldsymbol{J}_{\lambda_{n},AA}^{\star}\bigr)^{-1}\boldsymbol{R}_{n}.$
@@
$s_{n}^{1/2}P_{A}\bigl\{\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\bigr\}\longrightarrow_{d}N\bigl(\boldsymbol{0},\boldsymbol{K}^{\star}_{AA}\bigr).$
The rate condition $s_{n}^{1/2}(\lambda_{n}-\lambda)\to 0$ implies
$s_{n}^{1/2}(\lambda_{n}-\lambda)\bigl(\boldsymbol{J}_{\lambda_{n},AA}^{\star}\bigr)^{-1}\nabla_{A}\rho(\boldsymbol{\theta}^{\lambda})=o_{p}(1).$
By Assumption D.1 the active-set penalized curvature $\boldsymbol{J}_{\lambda,AA}^{\star}$ is nonsingular, and by continuity $\boldsymbol{J}_{\lambda_{n},AA}^{\star}\to\boldsymbol{J}_{\lambda,AA}^{\star}$, so that $\bigl(\boldsymbol{J}_{\lambda_{n},AA}^{\star}\bigr)^{-1}$ remains bounded in probability. Together with (26), the estimating equation above yields
$s_{n}^{1/2}\bigl\|\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}\bigr\|=O_{p}(1)$
by a standard $Z$-estimation rate argument as shown in vandervaart1998asymptotic. In particular, (26) implies
$s_{n}^{1/2}\|\boldsymbol{R}_{n}\|=o_{p}(1).$ (27)
Hence
$s_{n}^{1/2}\bigl\|\bigl(\boldsymbol{J}_{\lambda_{n},AA}^{\star}\bigr)^{-1}\boldsymbol{R}_{n}\bigr\|=o_{p}(1).$
Therefore,
$s_{n}^{1/2}\bigl(\hat{\boldsymbol{\theta}}_{n,A}^{\mathrm{pen}}-\boldsymbol{\theta}^{\lambda}_{A}\bigr)=-\bigl(\boldsymbol{J}_{\lambda_{n},AA}^{\star}\bigr)^{-1}s_{n}^{1/2}P_{A}\bigl\{\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\bigr\}+o_{p}(1).$
Since $\boldsymbol{J}_{\lambda_{n},AA}^{\star}\to\boldsymbol{J}_{\lambda,AA}^{\star}$, Slutsky’s theorem yields the claimed normal limit with covariance $(\boldsymbol{J}_{\lambda,AA}^{\star})^{-1}\boldsymbol{K}^{\star}_{AA}(\boldsymbol{J}_{\lambda,AA}^{\star})^{-1}$. ∎
Assumption D.1 and Lemma D.3 show that, under the non-smooth setting, the active coordinates behave as in the smooth case, with curvature matrix $\boldsymbol{J}_{\lambda,AA}^{\star}$ and variability $\boldsymbol{K}_{AA}^{\star}$. The proofs of Proposition 3.1, Theorem 3.2, Lemma 4.1, Lemma 4.2, and Proposition 4.3 use only local quadratic expansions and the central limit theorem. Repeating those arguments with all matrices and vectors restricted to $A$ yields the same Gaussian limits on the active coordinates after replacing $\boldsymbol{J}^{\star}$, $\boldsymbol{J}_{\lambda}^{\star}$, $\boldsymbol{K}^{\star}$, and $\boldsymbol{V}_{\mathrm{target}}^{\star}$ by their $AA$-blocks. Coordinates in $A^{c}$ may exhibit boundary phenomena, and their asymptotic distribution need not be Gaussian. Accordingly, our asymptotic statements and calibrated inference for non-smooth penalties are reported conditionally on the active set $A$.